# Transport alpha divergences

### v2

@@ 3.3. Properties
In this section, we show that there are several dualities and convexity properties for transport alpha divergences. The proofs are based on the fact that transport alpha divergences are generalized Bregman divergences in Wasserstein-$2$ space.
@@
    $\mathrm{D}_{\mathrm{T},\alpha}(p\|q)=\mathrm{D}_{\mathrm{T},-\alpha}(q|p).$
@@
(ii) The duality is from equation (7). For any $z_{1}$, $z_{2}>0$, we have $f_{\mathrm{T},\alpha}(\frac{z_{1}}{z_{2}})=f_{\mathrm{T},-\alpha}(\frac{z_{2}}{z_{1}})$. Replacing $z_{1}$, $z_{2}$ by QDFs $Q^{\prime}_{p}$, $Q^{\prime}_{q}$, respectively, and using (7), we finish the proof.
Proposition 3 (Taylor expansions in Wasserstein-$2$ spaces).
@@
$\begin{split}\mathrm{D}_{\mathrm{T},\alpha}(p\|q)=&\frac{1}{2}\int_{0}^{1}|\frac{Q^{\prime}_{p}(u)-Q^{\prime}_{q}(u)}{Q^{\prime}_{q}(u)}|^{2}du+\frac{\alpha-3}{6}\int_{0}^{1}\Big{(}\frac{Q^{\prime}_{p}(u)-Q^{\prime}_{q}(u)}{Q^{\prime}_{q}(u)}\Big{)}^{3}du\\ &+{O(\int_{0}^{1}|\frac{Q^{\prime}_{p}(u)-Q^{\prime}_{q}(u)}{Q^{\prime}_{q}(u)}|^{4}du)}.\end{split}$
@@
$f_{\mathrm{T},\alpha}(\frac{Q^{\prime}_{p}(u)}{Q^{\prime}_{q}(u)})=f_{\mathrm{T},\alpha}(1+h(u)),$
@@
We next represent transport alpha divergences in terms of generalized Bregman divergences in Wasserstein-$2$ spaces. Denote a function $\mathrm{D}_{-1}=\mathrm{D}_{\mathrm{IS}}\colon\mathbb{R}_{+}^{2}\rightarrow\mathbb{R}_{+}$, such that for $z_{1}$, $z_{2}\in\mathbb{R}_{+}$,
@@
Here, the notation $\mathrm{D}_{-1}$ is a (1)(-1)-divergence, which is a Bregman divergence with a potential function
@@
Here, $\mathrm{D}_{-1}$ is the Itakura–Saito divergence for the point divergence function.
Theorem 1 (Alpha–Itakura–Saito divergences in Wasserstein-$2$ spaces).
@@
$\begin{split}\mathrm{D}_{\mathrm{T},\alpha}(p\|q)=&\frac{1}{\alpha^{2}}\int_{0}^{1}\mathrm{D}_{-1}\big{(}Q_{p}^{\prime}(u)^{\alpha}|Q_{q}^{\prime}(u)^{\alpha}\big{)}du.\end{split}$
@@
$\begin{split}\mathrm{D}_{\mathrm{T},\alpha}(p\|q)=&\frac{1}{\alpha^{2}}\int_{0}^{1}\Big{[}\Psi(Q_{p}^{\prime}(u)^{\alpha})-\Psi(Q^{\prime}_{q}(u)^{\alpha})-\Psi^{\prime}(Q^{\prime}_{q}(u)^{\alpha})\cdot(Q^{\prime}_{p}(u)^{\alpha}-Q^{\prime}_{q}(u)^{\alpha})\Big{]}du.\end{split}$ (8)
@@
$\mathrm{D}_{\mathrm{T},\alpha}(p\|q)=\int_{0}^{1}f_{\mathrm{T},\alpha}(\frac{Q^{\prime}_{p}(u)}{Q^{\prime}_{q}(u)})du=\frac{1}{\alpha^{2}}\int_{0}^{1}{\mathrm{D}_{-1}\big{(}Q_{p}^{\prime}(u)^{\alpha}|Q_{q}^{\prime}(u)^{\alpha}\big{)}}du.$
From the fact of $\mathrm{D}_{-1}$ is a Bregman divergence function, we have
@@
$\begin{split}\int_{0}^{1}\Psi(Q^{\prime}_{q}(u)^{\alpha})du=&-\alpha\int_{0}^{1}\log(\frac{d}{du}Q^{\prime}_{q}(u))du\\ =&-\alpha\int_{0}^{1}\log\frac{1}{q(x)}q(x)dx=\alpha\int_{\Omega}q(x)\log q(x)dx.\end{split}$
@@
Corollary 2 (Generalized Pythagorean theorem in Wasserstein-$2$ spaces).
@@
$\left\{\begin{aligned} &\frac{1}{\alpha^{2}}\int_{0}^{1}\big{(}Q^{\prime}_{p}(u)^{\alpha}-Q^{\prime}_{q}(u)^{\alpha}\big{)}\cdot\big{(}\frac{1}{Q^{\prime}_{r}(u)^{\alpha}}-\frac{1}{Q^{\prime}_{q}(u)^{\alpha}}\big{)}du=0,&\quad\mathrm{if}\quad\alpha\neq 0;\\ &\int_{0}^{1}\log\frac{Q^{\prime}_{p}(u)}{Q^{\prime}_{q}(u)}\cdot\log\frac{Q^{\prime}_{r}(u)}{Q^{\prime}_{q}(u)}du=0,&\quad\mathrm{if}\quad\alpha\neq 0.\end{aligned}\right$. (11)
@@
$\mathrm{D}_{\mathrm{T},\alpha}(p\|q)+\mathrm{D}_{\mathrm{T},\alpha}(q|r)=\mathrm{D}_{\mathrm{T},\alpha}(p|r).$
@@
$\begin{split}&\mathrm{D}_{\mathrm{T},\alpha}(p\|q)+\mathrm{D}_{\mathrm{T},\alpha}(q|r)\\ =&\frac{1}{\alpha^{2}}\int_{0}^{1}\Big{[}\Psi(K_{p})+\Psi^{*}(K_{q}^{*})-K^{*}_{q}\cdot K_{p}+\Psi(K_{q})+\Psi^{*}(K_{r}^{*})-K^{*}_{r}\cdot K_{q}\Big{]}du\\ =&\frac{1}{\alpha^{2}}\int_{0}^{1}\Big{[}\Psi(K_{p})+\Psi^{*}(K_{r}^{*})-K_{p}\cdot K^{*}_{r}+K_{p}\cdot K^{*}_{r}+K_{q}\cdot K_{q}^{*}-K^{*}_{q}\cdot K_{p}-K^{*}_{r}\cdot K_{q}\Big{]}du\\ =&\mathrm{D}_{\mathrm{T},\alpha}(p|r)+\frac{1}{\alpha^{2}}\int_{0}^{1}(K_{p}-K_{q})\cdot(K^{*}_{r}-K^{*}_{q})du.\end{split}$
From the orthogonal condition (11), we finish the proof for $\alpha\neq 0$. For $\alpha=0$, the proof is from the fact that we use the coordinate $\hat{K}_{p}=\log Q^{\prime}_{p}(u)$, under which the transport alpha divergence is an Euclidean distance. The Pythagorean theorem is simple to show. ∎
We can also present the orthogonal condition (11) in terms of pushforward mapping functions.
@@
We remark that the proposed divergences are canonical divergences [6] associated with the transport Hessian metric. In the appendix, we first review the transprot Hessian meric, the transport KL divergence, and then provide the proof of canonical divergence.
@@ 3.4. Transport alpha geodesic
Given two probability density functions $p$, $q\in\mathcal{P}(\Omega)$ and $\alpha\in\mathbb{R}$, the transport alpha geodesic is defined as below. Denote a mapping function $T_{\alpha}\colon[0,1]\times\Omega\rightarrow\Omega$. Consider a one-parameter family of partial differential equations:
$\partial_{tt}\partial_{x}T_{\alpha}(t,x)-(\alpha+1)\frac{(\partial_{t}\partial_{x}T_{\alpha}(t,x))^{2}}{\partial_{x}T_{\alpha}(t,x)}=0,$ (12)
@@
Let $T$ be defined in (6). Assume $T^{\prime}(x)\neq 0$ for all $x\in\Omega$. A solution of transport alpha geodesic is given below. The mapping function $T_{\alpha}$ satisfies
@@
Equivalently, denote $r_{\alpha}(t,\cdot)=T_{\alpha}(t,\cdot)_{\#}q$, and write $Q_{r_{\alpha}}(t,\cdot)$, $\partial_{u}Q_{r_{\alpha}}(t,u)$ as the quantile function, quantile density function of probability density $r_{\alpha}(t,\cdot)$, respectively. Then the transport alpha geodesics in QDFs satisfies
@@
$\partial_{tt}(\partial_{x}T_{\alpha}(t,x))^{-\alpha}=0,$
@@
$\begin{split}(\partial_{x}T_{\alpha}(t,x))^{-\alpha}=&t(\partial_{x}T_{\alpha}(1,x))^{-\alpha}+(1-t)(\partial_{x}T_{\alpha}(0,x))^{-\alpha}\\ =&tT^{\prime}(x)^{-\alpha}+(1-t).\end{split}$
This finishes the first part of the proof. By changing the variable $u=F_{q}(x)$, we have
@@
$\partial_{u}Q_{r_{\alpha}}(t,u)=\Big{(}(1-t)+t(\frac{Q^{\prime}_{p}(u)}{Q^{\prime}_{q}(u)})^{-\alpha}\Big{)}^{-\frac{1}{\alpha}}\cdot Q^{\prime}_{q}(u).$
@@
Proposition 4 can be explained as follows. If $\alpha=-1$, transport-(1)(-1) geodesic also satisfies the geodesic equation in Wasserstein-$2$ space, which is “transportation flat”, meaning that the flatness in the pushforward mapping functions:
@@
While, if $\alpha=1$, the transport-$1$ geodesic is an “inverse Jacobi transportation flat” curve. The mapping function pushforwards the density $q$ to $p$ flatly from the following equation:
@@
If $\alpha=0$, the transport-0 geodesic is a geodesic equation in the transport Hessian metric of negative Boltzmann-Shannon entropy [18, 19]. From now on, we call (13) the $m$-geodesic in Wasserstein-$2$ space, while name (14) the $e$-geodesic in Wasserstein-$2$ space.
@@
Consider two Gaussian distributions $p=\mathcal{N}(0,\Sigma_{p})$, $q=\mathcal{N}(0,\Sigma_{q})$, where $\Sigma_{p}$, $\Sigma_{q}>0$ are covariances of $p$, $q$, respectively. Then $T(x)=\Sigma_{p}^{\frac{1}{2}}\Sigma_{q}^{-\frac{1}{2}}x$. Thus,
$\partial_{x}T_{-1}(t,x)=(1-t)+t\Sigma_{p}^{\frac{1}{2}}\Sigma_{q}^{-\frac{1}{2}},\quad\partial_{x}T_{0}(t,x)=(\Sigma_{p}^{\frac{1}{2}}\Sigma_{q}^{-\frac{1}{2}})^{t},\quad\partial_{x}T_{1}(t,x)=\frac{1}{(1-t)+t\Sigma_{p}^{-\frac{1}{2}}\Sigma_{q}^{\frac{1}{2}}}.$
The following figure demonstrates the above three alpha geodesics.
Figure 1. Three curves for $\partial_{x}T_{\alpha}(t,x)$ with $\Sigma_{p}=1$, $\Sigma_{q}=4$. Red: $\alpha=1$. Black: $\alpha=0$. Blue $\alpha=-1$.

### v3

@@ 3.3. Properties
In this section, we show that there are several dualities and convexity properties for transport alpha divergences. These proofs are based on the facts that transport alpha divergences are generalized Bregman divergences in Wasserstein-$2$ space.
@@
    $\mathrm{D}_{\mathrm{T},\alpha}(p\|q)=\mathrm{D}_{\mathrm{T},-\alpha}(q\|p).$
@@
(ii) The duality is from equation (7). For any $z_{1}$, $z_{2}>0$, we have $f_{\mathrm{T},\alpha}(\frac{z_{1}}{z_{2}})=f_{\mathrm{T},-\alpha}(\frac{z_{2}}{z_{1}})$. That is
$\begin{split}f_{\mathrm{T},\alpha}\left(\frac{z_{1}}{z_{2}}\right)=&\left\{\begin{aligned} &\frac{1}{\alpha^{2}}\left((\frac{z_{1}}{z_{2}})^{\alpha}-\alpha\log\frac{z_{1}}{z_{2}}-1\right),&\quad\mathrm{if}\quad\alpha\neq 0;\\ &\frac{1}{2}|\log\frac{z_{1}}{z_{2}}|^{2},&\quad\mathrm{if}\quad\alpha=0,\end{aligned}\right.\\ =&\left\{\begin{aligned} &\frac{1}{\alpha^{2}}\left((\frac{z_{2}}{z_{1}})^{-\alpha}-(-\alpha)\log\frac{z_{2}}{z_{1}}-1\right),&\quad\mathrm{if}\quad\alpha\neq 0;\\ &\frac{1}{2}\left|\log\frac{z_{2}}{z_{1}}\right|^{2},&\quad\mathrm{if}\quad\alpha=0,\end{aligned}\right.\\ =&f_{\mathrm{T},-\alpha}(\frac{z_{2}}{z_{1}}).\end{split}$
Replacing $z_{1}$, $z_{2}$ by QDFs $Q^{\prime}_{p}$, $Q^{\prime}_{q}$, respectively, and using (7), we finish the proof.
Proposition 3 (Taylor expansions in Wasserstein-$2$ space).
@@
$\begin{split}\mathrm{D}_{\mathrm{T},\alpha}(p\|q)=&\frac{1}{2}\int_{0}^{1}\left|\frac{Q^{\prime}_{p}(u)-Q^{\prime}_{q}(u)}{Q^{\prime}_{q}(u)}\right|^{2}du+\frac{\alpha-3}{6}\int_{0}^{1}\Big(\frac{Q^{\prime}_{p}(u)-Q^{\prime}_{q}(u)}{Q^{\prime}_{q}(u)}\Big)^{3}du\\ &+{O\left(\int_{0}^{1}|\frac{Q^{\prime}_{p}(u)-Q^{\prime}_{q}(u)}{Q^{\prime}_{q}(u)}|^{4}du\right)}.\end{split}$
@@
$f_{\mathrm{T},\alpha}\left(\frac{Q^{\prime}_{p}(u)}{Q^{\prime}_{q}(u)}\right)=f_{\mathrm{T},\alpha}(1+h(u)),$
@@
We next represent transport alpha divergences in terms of generalized Bregman divergences in Wasserstein-$2$ space. Denote a function $\mathrm{D}_{-1}=\mathrm{D}_{\mathrm{IS}}\colon\mathbb{R}_{+}^{2}\rightarrow\mathbb{R}_{+}$, such that for $z_{1}$, $z_{2}\in\mathbb{R}_{+}$,
@@
The notation $\mathrm{D}_{-1}$ is a (1)(-1)-divergence, which is a Bregman divergence with a potential function
@@
Here $\mathrm{D}_{-1}$ is the Itakura–Saito divergence on $\mathbb{R}_{+}$.
Theorem 1 (Alpha–Itakura–Saito divergences in Wasserstein-$2$ space).
@@
$\begin{split}\mathrm{D}_{\mathrm{T},\alpha}(p\|q)=&\frac{1}{\alpha^{2}}\int_{0}^{1}\mathrm{D}_{-1}\left(Q_{p}^{\prime}(u)^{\alpha}|Q_{q}^{\prime}(u)^{\alpha}\right)du.\end{split}$
@@
$\begin{split}\mathrm{D}_{\mathrm{T},\alpha}(p\|q)=&\frac{1}{\alpha^{2}}\int_{0}^{1}\left[\Psi(Q_{p}^{\prime}(u)^{\alpha})-\Psi\left(Q^{\prime}_{q}(u)^{\alpha}\right)-\Psi^{\prime}(Q^{\prime}_{q}(u)^{\alpha})\cdot\left(Q^{\prime}_{p}(u)^{\alpha}-Q^{\prime}_{q}(u)^{\alpha}\right)\right]du.\end{split}$ (8)
@@
$\mathrm{D}_{\mathrm{T},\alpha}(p\|q)=\int_{0}^{1}f_{\mathrm{T},\alpha}\left(\frac{Q^{\prime}_{p}(u)}{Q^{\prime}_{q}(u)}\right)du=\frac{1}{\alpha^{2}}\int_{0}^{1}{\mathrm{D}_{-1}\left(Q_{p}^{\prime}(u)^{\alpha}|Q_{q}^{\prime}(u)^{\alpha}\right)}du.$
From the fact that $\mathrm{D}_{-1}$ is a Bregman divergence function, we have
@@
$\begin{split}\int_{0}^{1}\Psi\left(Q^{\prime}_{q}(u)^{\alpha}\right)du=&-\alpha\int_{0}^{1}\log\left(\frac{d}{du}Q^{\prime}_{q}(u)\right)du\\ =&-\alpha\int_{0}^{1}\log\frac{1}{q(x)}q(x)dx=\alpha\int_{\Omega}q(x)\log q(x)dx.\end{split}$
@@
Corollary 2 (Generalized Pythagorean theorem in Wasserstein-$2$ space).
@@
$\left\{\begin{aligned} &\frac{1}{\alpha^{2}}\int_{0}^{1}\big(Q^{\prime}_{p}(u)^{\alpha}-Q^{\prime}_{q}(u)^{\alpha}\big)\cdot\left(\frac{1}{Q^{\prime}_{r}(u)^{\alpha}}-\frac{1}{Q^{\prime}_{q}(u)^{\alpha}}\right)du=0,&\quad\mathrm{if}\quad\alpha\neq 0;\\ &\int_{0}^{1}\log\frac{Q^{\prime}_{p}(u)}{Q^{\prime}_{q}(u)}\cdot\log\frac{Q^{\prime}_{r}(u)}{Q^{\prime}_{q}(u)}du=0,&\quad\mathrm{if}\quad\alpha\neq 0.\end{aligned}\right$. (11)
@@
$\mathrm{D}_{\mathrm{T},\alpha}(p\|q)+\mathrm{D}_{\mathrm{T},\alpha}(q\|r)=\mathrm{D}_{\mathrm{T},\alpha}(p\|r).$
@@
$\begin{split}&\mathrm{D}_{\mathrm{T},\alpha}(p\|q)+\mathrm{D}_{\mathrm{T},\alpha}(q\|r)\\ =&\frac{1}{\alpha^{2}}\int_{0}^{1}\Big[\Psi(K_{p})+\Psi^{*}(K_{q}^{*})-K^{*}_{q}\cdot K_{p}+\Psi(K_{q})+\Psi^{*}(K_{r}^{*})-K^{*}_{r}\cdot K_{q}\Big]du\\ =&\frac{1}{\alpha^{2}}\int_{0}^{1}\Big[\Psi(K_{p})+\Psi^{*}(K_{r}^{*})-K_{p}\cdot K^{*}_{r}+K_{p}\cdot K^{*}_{r}+K_{q}\cdot K_{q}^{*}-K^{*}_{q}\cdot K_{p}-K^{*}_{r}\cdot K_{q}\Big]du\\ =&\mathrm{D}_{\mathrm{T},\alpha}(p\|r)+\frac{1}{\alpha^{2}}\int_{0}^{1}(K_{p}-K_{q})\cdot(K^{*}_{r}-K^{*}_{q})du.\end{split}$
From the orthogonal condition (11), we finish the proof for $\alpha\neq 0$. For $\alpha=0$, the proof is from the fact that we use the coordinate $\log Q^{\prime}_{p}(u)$, under which the transport alpha divergence is an Euclidean distance. The result is easy to show. This finishes the proof. ∎
We also present the orthogonal condition (11) in terms of transport maps.
@@
We remark that the proposed divergences are canonical divergences [6] associated with the transport Hessian metric. In the appendix, we first review the transport Hessian distance, the transport KL divergence, and then provide a proof of the canonical divergence for the transport alpha divergence.
Remark 3.
We also point out the other derivation of transport alpha divergences. We work on the potential function $\Psi(z)=-\log z$, $z>0$, and derive the alpha divergence of $\Psi(z)$ on $\mathbb{R}_{+}$. We then substitute the variable $z$ by the quantile density function in the integral of the domain $[0,1]$. This procedure derives the proposed transport alpha divergence, which is a one family generalization of the transport Hessian distance and the transport KL divergence.
@@ 3.4. Transport alpha geodesic
Given two probability density functions $p$, $q\in\mathcal{P}(\Omega)$ and $\alpha\in\mathbb{R}$, the transport alpha geodesic is defined as below. Denote a transport map function $T_{\alpha}\colon[0,1]\times\Omega\rightarrow\Omega$. Consider a one-parameter family of partial differential equations:
$\partial_{tt}\partial_{x}T_{\alpha}(t,x)-(\alpha+1)\frac{\big(\partial_{t}\partial_{x}T_{\alpha}(t,x)\big)^{2}}{\partial_{x}T_{\alpha}(t,x)}=0,$ (12)
@@
Let $T$ be defined in (6). Assume $T^{\prime}(x)\neq 0$ for all $x\in\Omega$. A solution of transport alpha geodesic is given below. The transport map function $T_{\alpha}$ satisfies
@@
Equivalently, denote $r_{\alpha}(t,\cdot)=T_{\alpha}(t,\cdot)_{\#}q$, and write $Q_{r_{\alpha}}(t,\cdot)$, $\partial_{u}Q_{r_{\alpha}}(t,u)$ as the quantile function, quantile density function of probability density function $r_{\alpha}(t,\cdot)$, respectively. Then the transport alpha geodesics in QDFs satisfies
@@
$\partial_{tt}\big(\partial_{x}T_{\alpha}(t,x)\big)^{-\alpha}=0,$
@@
$\begin{split}\big(\partial_{x}T_{\alpha}(t,x)\big)^{-\alpha}=&t(\partial_{x}T_{\alpha}(1,x))^{-\alpha}+(1-t)(\partial_{x}T_{\alpha}(0,x))^{-\alpha}\\ =&tT^{\prime}(x)^{-\alpha}+(1-t).\end{split}$
This finishes the first part of the proof.
By changing the variable $u=F_{q}(x)$, we have
@@
$\partial_{u}Q_{r_{\alpha}}(t,u)=\Big((1-t)+t\left(\frac{Q^{\prime}_{p}(u)}{Q^{\prime}_{q}(u)}\right)^{-\alpha}\Big)^{-\frac{1}{\alpha}}\cdot Q^{\prime}_{q}(u).$
@@
Proposition 4 can be explained as follows. If $\alpha=-1$, transport (1)(-1) geodesic also satisfies the geodesic equation in Wasserstein-$2$ space, which is “transportation flat”, meaning that the flatness in the transport maps, known as the McCann’s displacement interpolation:
@@
While, if $\alpha=1$, the transport $1$ geodesic is an “inverse Jacobi transportation flat” curve. The transport map function pushforwards the density $q$ to $p$ flatly from the following equation:
@@
If $\alpha=0$, the transport-0 geodesic is a geodesic equation in the transport Hessian metric of negative Boltzmann-Shannon entropy [19, 20]. From now on, we call (13) the mm–geodesic in Wasserstein-$2$ space, while name (14) the ee–geodesic in Wasserstein-$2$ space.
@@
Consider two Gaussian distributions $p=\mathcal{N}(0,\sigma_{p}^{2})$, $q=\mathcal{N}(0,\sigma_{q}^{2})$, where $\sigma_{p}$, $\sigma_{q}>0$ are standard variances of $p$, $q$, respectively. Then $T(x)=\sigma_{p}\sigma_{q}^{-1}x$. Thus, $\partial_{x}T_{\alpha}(t,x)=\Big((1-t)+t(\sigma_{p}\sigma_{q}^{-1})^{-\alpha}\Big)^{-\frac{1}{\alpha}}$. In this case, the transport alpha geodesic $(T_{\alpha}(t,\cdot))_{\#}q=r_{\alpha}(t,\cdot)$ satisfies a time-dependent Gaussian distribution, whose standard variance $\sigma_{\alpha}$ satisfies $\sigma_{\alpha}(t)=\sigma_{q}\cdot\partial_{x}T_{\alpha}(t,x)$. Hence, let $\alpha=-1,0,1$, we have
$\partial_{x}T_{-1}(t,x)=(1-t)+t\sigma_{p}\sigma_{q}^{-1},\quad\partial_{x}T_{0}(t,x)=(\sigma_{p}\sigma_{q}^{-1})^{t},\quad\partial_{x}T_{1}(t,x)=\frac{1}{(1-t)+t\sigma_{p}^{-1}\sigma_{q}}.$
And
$\sigma_{-1}(t)=(1-t)\sigma_{q}+t\sigma_{p},\quad\sigma_{0}(x)=\sigma_{p}^{t}\sigma_{q}^{1-t},\quad\sigma_{1}(t)=\frac{1}{(1-t)\sigma_{q}^{-1}+t\sigma_{p}^{-1}}.$
The following two figures demonstrate the above three transport alpha geodesics.
Figure 1. Three curves for $\partial_{x}T_{\alpha}(t,x)$ (left), and $\sigma_{\alpha}(t)$ (right) with $\sigma_{p}=1$, $\sigma_{q}=5$. Red: $\alpha=1$. Black: $\alpha=0$. Blue $\alpha=-1$.
We note that Figure 1 represents three geodesics in terms of the derivatives of transport map functions and standard variances, which are up to a constant ratio. For standard variances, we show that the transport (1)(-1) geodesic is a line interpolation between $\sigma_{p}$ and $\sigma_{q}$, the transport $1$ geodesic is the inverse function of the time variable, while the transport 0 geodesic is the power function of the time variable.
Remark 4.
We informally present the dualistic geometry for transport alpha geodesics. For simplicity of discussion, we study alpha-connections for quantile density functions. Denote a manifold $\mathcal{M}=\Big\{Q^{\prime}_{p}\in C^{\infty}[0,1]\colon Q_{p}^{\prime}(u)>0,\quad\textrm{for all $u\in[0,1]$}\Big\}$. Denote the tangent vector in the tangent space at $Q^{\prime}_{p}\in\mathcal{M}$ as $\mathcal{T}_{Q^{\prime}_{p}}\mathcal{M}=\Big\{\dot{Q}^{\prime}_{p}\in C^{\infty}[0,1]\Big\}$. Denote the alpha–connection $\mathbf{\Gamma}^{\alpha}\colon\mathcal{M}\times C^{\infty}[0,1]\times C^{\infty}[0,1]\rightarrow C^{\infty}[0,1]$ at the point $Q_{p}^{\prime}$ as
$\mathbf{\Gamma}^{\alpha}:=\mathbf{\Gamma}^{\alpha}(Q^{\prime}_{p},\dot{Q}^{\prime}_{p},\dot{Q}^{\prime}_{p})=-(\alpha+1)\frac{|\dot{Q}^{\prime}_{p}(u)|^{2}}{Q^{\prime}_{p}(u)}.$
Denote the 0–connection as $\mathbf{\Gamma}^{0}$, which is the Levi-Civita connection of the transport Hessian metric:
$\mathbf{\Gamma}^{0}:=\mathbf{\Gamma}^{0}(Q^{\prime}_{p},\dot{Q}^{\prime}_{p},\dot{Q}^{\prime}_{p})=-\frac{|\dot{Q}^{\prime}_{p}(u)|^{2}}{Q^{\prime}_{p}(u)}.$
Denote the 11–connection as $\mathbf{\Gamma}$, which equals to
$\mathbf{\Gamma}:=\mathbf{\Gamma}^{1}(Q^{\prime}_{p},\dot{Q}^{\prime}_{p},\dot{Q}^{\prime}_{p})=-2\frac{|\dot{Q}^{\prime}_{p}(u)|^{2}}{Q^{\prime}_{p}(u)}.$
Denote the (1)(-1)–connection as $\mathbf{\Gamma}^{*}$, such that
$\mathbf{\Gamma}^{*}:=\mathbf{\Gamma}^{-1}(Q^{\prime}_{p},\dot{Q}^{\prime}_{p},\dot{Q}^{\prime}_{p})=0.$
The following condition holds: for any $\alpha\in\mathbb{R}$, we have
$\mathbf{\Gamma}^{\alpha}=\frac{1+\alpha}{2}\mathbf{\Gamma}+\frac{1-\alpha}{2}\mathbf{\Gamma}^{*}.$
If $\alpha=0$, then $\mathbf{\Gamma}^{0}=\frac{1}{2}(\mathbf{\Gamma}+\mathbf{\Gamma}^{*})$. If we further denote an operator $\mathbf{C}\colon\mathcal{M}\times C^{\infty}[0,1]\times C^{\infty}[0,1]\rightarrow C^{\infty}[0,1]$, such that
$\mathbf{C}:=\mathbf{C}(Q^{\prime}_{p},\dot{Q}^{\prime}_{p},\dot{Q}^{\prime}_{p})=2\cdot\frac{|\dot{Q}^{\prime}_{p}(u)|^{2}}{Q^{\prime}_{p}(u)},$
then
$\mathbf{\Gamma}^{\alpha}=\mathbf{\Gamma}^{0}-\frac{\alpha}{2}\mathbf{C},\quad\mathbf{\Gamma}^{-\alpha}=\mathbf{\Gamma}^{0}+\frac{\alpha}{2}\mathbf{C}.$
Remark 5.
We remark a relation between the geodesic and the orthogonal condition in Corollary 2. For $\alpha\neq 0$, note that for $\gamma_{\alpha}$ connecting $p$, $q$ in Definition 2, we have
$\partial_{t}\Big(\partial_{u}Q_{r_{\alpha}}(t,u)\Big)^{-\alpha}=Q^{\prime}_{p}(u)^{-\alpha}-Q^{\prime}_{q}(u)^{-\alpha}.$
Suppose that $l_{-\alpha,p,q}(t,\cdot)$, $t\in[0,1]$ is the transport negative alpha geodesic connecting $p$, $q$, such that $l_{-\alpha,p,q}(0,\cdot)=p(\cdot)$, $l_{-\alpha,p,q}(1,\cdot)=q(\cdot)$. And $l_{\alpha,r,q}(t,\cdot)$, $t\in[0,1]$, is the transport alpha geodesic connecting $r$, and $q$, such that $l_{\alpha,r,q}(0,\cdot)=r(\cdot)$, $l_{\alpha,r,q}(1)=q(\cdot)$. Then the orthogonal condition in Corollary 2 can be written as
$\int_{0}^{1}\partial_{t}\Big(\partial_{u}Q_{l_{-\alpha,p,q}}(t,u)\Big)^{\alpha}\cdot\partial_{t}\Big(\partial_{u}Q_{l_{\alpha,r,q}}(t,u)\Big)^{-\alpha}du=0.$
Similar results also hold for $\alpha=0$.